An oil exploration firm is to drill 9 wells, with each having probability 0.15 of producing oil.
(Express your answers as fractions in simplest form or as decimals rounded to four decimal places.)
a. Find the expected number of wells that would produce oil.
b. Find the standard deviation of the number wells that would produce oil.

Answer by stanbon(75887)

An oil exploration firm is to drill 9 wells, with each having probability 0.15 of producing oil.
(Express your answers as fractions in simplest form or as decimals rounded to four decimal places.)
a. Find the expected number of wells that would produce oil.
Expected value = mean = n*p = 9*0.15 = 1.35
--------------------------
b. Find the standard deviation of the number wells that would produce oil.
std = sqrt(npq) = sqrt(1.35*0.85) = 1.07
-------------
Cheers,
Stan H.
-------------